5/X = X/20
Do I use the positive solution, the negative solution, or both?

Found 2 solutions by checkley77, solver91311:
Answer by checkley77(12844)

5/X = X/20
X^2=100
X=SQRT100
X=+/-10 ANS.
BOTH = & - ANSWERS ARE VALID.

Answer by solver91311(24713)



Well, since:



and



From a purely mathematical point of view, either solution makes the original equation a true statement, and therefore both solutions are members of the solution set.

On the other hand, if you are using a proportion to solve a physical problem where the answer is a positive measure of something, then you would certainly exclude the negative root. On the other hand, you might have a problem that requires you to find a root that is less than zero -- hence you would exclude the positive root.

So, you tell me.


John